OpenStudy (anonymous):

a car speeds up from 15 m/s to 30 m/s in 12 seconds. what is the acceleration of the car

OpenStudy (anonymous):

\[a=\Delta v/\Delta t\] if continuous. 30-15/12=15/12=1.25m/s^2